# Optimization tools for distance-preserving flag fault-tolerant error correction

### v3

@@ IV.2.1 Counting faults in correlated and uncorrelated bit histories
In this work, we further extend the fault counting idea to flag FTQEC in which flag circuits with single flag qubit are used for syndrome extraction. Below, we will discuss two types of adaptive time decoders with different stop conditions, namely one-tailed and two-tailed adaptive time decoders. Both protocols are applicable to any stabilizer code as long as flag circuits for the code that give a distinguishable fault set can be found. The flag FTQEC protocol with one-tailed adaptive time decoder satisfies the FTQEC conditions in Definition 4, thus it is applicable to any fault-tolerant quantum computation as long as the fault-tolerant implementation of other operations (gate, state preparation, or measurement) also satisfies the revised definition of fault tolerance which considers the number of faults instead of the weight of the error [35]. Meanwhile, the flag FTQEC protocol with two-tailed adaptive time decoder does not satisfy the FTQEC conditions in Definition 4 as the output error may correspond to a nontrivial cumulative flag vector, hence it is only applicable to quantum memory. Nevertheless, for a self-dual CSS code, the FTQEC protocol with the two-tailed adaptive time decoder can be applied to any fault-tolerant Clifford computation if the cumulative flag vector is processed appropriately. An analysis of this extension will be discussed in Section IV.3.
@@ IV.3 Extended techniques for CSS codes
Table 2: A list of required classical processing operations on the remaining cumulative flag vector in case that a logical Clifford gate is performed between two FTQEC routines. With these operations, a flag FTQEC protocol with two-tailed adaptive time decoder or separated $X$ and $Z$ counting is applicable to any fault-tolerant Clifford computation, given that the CSS code is self-dual.
@@ IV.3.2 Classical processing of the remaining cumulative flag vector
Nevertheless, for any self-dual CSS code, a flag FTQEC protocol with two-tailed adaptive time decoder or separated $X$ and $Z$ counting (or both) can made applicable to any fault-tolerant Clifford computation. For example, let us consider an application of a logical Hadamard gate $\bar{H}$ between two FTQEC routines. Suppose that the first FTQEC routine causes an output error $E_{x}\cdot E_{z}$ and the remaining cumulative flag vector is $(\vec{F}_{x},\vec{F}_{z})$. Without a logical Hadamard gate, $E_{x}$ and $E_{z}$ can be corrected using $\vec{F}_{z}$ and $\vec{F}_{x}$, respectively. A logical Hadamard gate transforms an $X$-type error to a $Z$-type error of the same form, and vice versa. Because the $X$-type and $Z$-type generators are of the same form, possible fault combinations for both types of errors are also of the same form. To correct the transformed error $\bar{H}(E_{x}\cdot E_{z})\bar{H}^{\dagger}$ in the second FTQEC routine, one needs to swap the $X$-type and $Z$-type cumulative flag vector; that is, the initial flag vector for the second FTQEC routine must be $(\vec{F}_{z},\vec{F}_{x})$.
We can apply similar ideas for flag information processing to logical $S$ and logical CNOT gates. The summary of the classical processing operations for logical $H$, $S$, and CNOT gates is provided in Table 2. Because $\{H,S,\text{CNOT}\}$ generates the Clifford group, a flag FTQEC protocol with two-tailed adaptive time decoder or separated $X$ and $Z$ counting is applicable to any fault-tolerant Clifford computation given that the CSS code is self-dual. Note that the magic state distillation and injection [56, 57] use only Clifford operations. Thus, our techniques are also applicable to fault-tolerant universal quantum computation given that high-fidelity magic states are provided.
@@ V.1 Methods
Our optimization tools for space and time decoders including the compact lookup table construction, the MIM technique, and the adaptive time decoders for flag FTQEC are applicable to any stabilizer code. However, we focus on a specific family of codes where the aforementioned tools can be simplified and extended techniques, including separated $X$ and $Z$ decoding and classical processing of flag information are applicable—the family of self-dual CSS codes in which the number of physical qubits is odd, the number of logical qubits is 1, and logical $X$ and $Z$ operators are transversal. To evaluate the performance of our tools, we simulate FTQEC protocols on the $[\![(3d^{2}+1)/4,1,d]\!]$ hexagonal color codes [41] of distance 3, 5, 7 and 9. These codes are planar topological codes with configurations displayed in Fig. 6. For each code, stabilizer generators are measured using the syndrome extraction circuits with single flag ancilla, as depicted in Fig. 4. It was proven that for the hexagonal code of any distance, using flag circuits of this form preserves the code distance regardless of the gate orderings [32, 35]). The simulation is implemented under the circuit-level depolarizing noise model specified in Section II.1. As there is no idling noise in our error model, the syndromes can be extracted sequentially.
@@
FIGURE 6: The studied members of the hexagonal color code, for distances 3, 5, 7, and 9 (right to left). Qubits are on the vertices and stabilizer generators are the plaquettes. As the codes are self-dual CSS codes, both the $X$ and $Z$ stabilizer generators are described by the same layout.
@@ V.2 The overall effect of optimization tools
We first compare two protocols: (1) the FTQEC protocol with Shor time decoder without the MIM technique (the protocol in which none of our optimization tools are applied) and (2) the FTQEC protocol with the MIM technique and the two-tailed adaptive time decoder with the ZX strategy (the best FTQEC protocol in this work which is compatible with any Clifford computation on a self-dual CSS code). The logical error rate $p_{L}$ vs physical error rate $p$ for hexagonal color codes of distance 3, 5, 7, and 9 are plotted in Fig. 7. Our results show that for each code, applying the optimization tools can significantly improve the pseudothreshold (the intersection between each plot and the $p_{L}=2p/3$ line). Furthermore, the optimized decoder yields orders of magnitude improvements in the logical error rate in the $p=10^{-4}$ error regime.
@@ VI Discussions and conclusions
In this work, we focus on flag FTQEC with lookup table decoding and improvements to a decoder consisting of a time decoder and a space decoder. For the space decoder, we first develop a technique to build the lookup table more efficiently in Section III.1. With our lookup table construction method, the lookup table for a self-dual CSS code requires at least 87.5% less memory compared to the lookup table for a generic stabilizer code. The construction method also verifies the distinguishability of the fault set corresponding to flag circuits for syndrome measurements. Our construction also leads to the notion of the fault code, a linear code corresponding to the faults under circuit-level noise, which simplifies the verification of the distance of the protocol. More efficient decoding schemes for the fault code can be an interesting avenue to explore in future work.
@@
For the time decoder, we generalize the adaptive syndrome measurement technique from the previous work [38] (which is applicable to Shor-style error correction [1]) to flag FTQEC, and develop one-tailed and two-tailed adaptive time decoders in Section IV.2. For a general stabilizer code in which flag FTQEC is possible, the one-tailed decoder is preferable as it is compatible with any fault-tolerant quantum computation, while the two-tailed decoder is applicable to quantum memory only. Nevertheless, for self-dual CSS codes, the two-tailed decoder is applicable to any fault-tolerant computation built from Clifford gates and application of $T$ gates by gate teleportation using high-fidelity magic states with the help of the classical processing technique on cumulative flag vectors developed in Section IV.3. The effect of the adaptive time decoders on the simulated storage is discussed in Section V.4. We observe that our adaptive time decoders can improve the pseudothresholds compared to the non-adaptive (Shor) time decoder while preserving the code distance. The two-tailed decoder also outperforms the one-tailed decoder.
@@
We can further improve the performance of adaptive time decoders on self-dual CSS codes by using the separated $X$ and $Z$ counting technique described in Section IV.3. Here, we estimate the number of faults occurred from the measurement of generators of the first type (either $X$ or $Z$) and then use that information in the measurement of generators of the second type. The effect of this technique can be found in Section V.5. When the logical $|0\rangle$ state is stored, we find that the protocol that measures $Z$-type generators before $X$-type generators performs the best. We see no significant difference in the protocol that measures $X$-type generators before $Z$-type generators, and the protocol that measures $X$-type and $Z$-type generators jointly. Thus, the separated $X$ and $Z$ counting provides an advantage only for certain input states depending on the measurement order.
@@ Appendix A Memory footprint savings in the lookup table
If the code is self-dual, then the two tables coincide, $T:=T_{X}=T_{Z}$, $r:=r_{X}=r_{Z}$. Thus,
$M_{CSS,CRO,SD}$ $=\frac{1}{2}\left(2(r_{Z}T_{Z}+r_{X}T_{X})+2n+k(T_{Z}+T_{X})\right)$
@@
which is consistent with having $T+1$ entries, with a map key of $n-k$ bits (with $(n-k)/2$ for flags and for generator bits) and a value with $k$ bits. If we would view the self-dual CSS code as a generic stabilizer code, we would get $T_{stab}=1+2T+T_{XZ}$ and thus the ratio is,
$M_{CSS,CRO,SD}/M_{stab}$ $=\frac{n(T+1)}{T_{stab}(4n-2k)}$
@@
where we used the fact that $T_{XZ}$ must be at least 1 for $t>0$ and a non-trivial encoding. This upper bound means that at a zero rate code, leveraging the structure of a self-dual CSS code and the CROs can create a memory footprint less than 12.5% that of the memory footprint of a lookup table if we viewed the code as a stabilizer code.

### v4

@@ IV.2.1 Counting faults in correlated and uncorrelated bit histories
In this work, we further extend the fault counting idea to flag FTQEC in which flag circuits with single flag qubit are used for syndrome extraction. Below, we will discuss two types of adaptive time decoders with different stop conditions, namely one-tailed and two-tailed adaptive time decoders. Both protocols are applicable to any stabilizer code as long as flag circuits for the code that give a distinguishable fault set can be found. The flag FTQEC protocol with one-tailed adaptive time decoder satisfies the FTQEC conditions in Definition 4, thus it is applicable to any fault-tolerant quantum computation as long as the fault-tolerant implementation of other operations (gate, state preparation, or measurement) also satisfies the revised definition of fault tolerance which considers the number of faults instead of the weight of the error [35]. Meanwhile, the flag FTQEC protocol with two-tailed adaptive time decoder does not satisfy the FTQEC conditions in Definition 4 as the output error may correspond to a nontrivial cumulative flag vector, hence it is only applicable to quantum memory. Nevertheless, for a self-orthogonal CSS code, the FTQEC protocol with the two-tailed adaptive time decoder can be applied to any fault-tolerant Clifford computation if the cumulative flag vector is processed appropriately. An analysis of this extension will be discussed in Section IV.3.
@@ IV.3 Extended techniques for CSS codes
Table 2: A list of required classical processing operations on the remaining cumulative flag vector in case that a logical Clifford gate is performed between two FTQEC routines. With these operations, a flag FTQEC protocol with two-tailed adaptive time decoder or separated $X$ and $Z$ counting is applicable to any fault-tolerant Clifford computation, given that the CSS code is self-orthogonal.
@@ IV.3.2 Classical processing of the remaining cumulative flag vector
Nevertheless, for any self-orthogonal CSS code, a flag FTQEC protocol with two-tailed adaptive time decoder or separated $X$ and $Z$ counting (or both) can made applicable to any fault-tolerant Clifford computation. For example, let us consider an application of a logical Hadamard gate $\bar{H}$ between two FTQEC routines. Suppose that the first FTQEC routine causes an output error $E_{x}\cdot E_{z}$ and the remaining cumulative flag vector is $(\vec{F}_{x},\vec{F}_{z})$. Without a logical Hadamard gate, $E_{x}$ and $E_{z}$ can be corrected using $\vec{F}_{z}$ and $\vec{F}_{x}$, respectively. A logical Hadamard gate transforms an $X$-type error to a $Z$-type error of the same form, and vice versa. Because the $X$-type and $Z$-type generators are of the same form, possible fault combinations for both types of errors are also of the same form. To correct the transformed error $\bar{H}(E_{x}\cdot E_{z})\bar{H}^{\dagger}$ in the second FTQEC routine, one needs to swap the $X$-type and $Z$-type cumulative flag vector; that is, the initial flag vector for the second FTQEC routine must be $(\vec{F}_{z},\vec{F}_{x})$.
We can apply similar ideas for flag information processing to logical $S$ and logical CNOT gates. The summary of the classical processing operations for logical $H$, $S$, and CNOT gates is provided in Table 2. Because $\{H,S,\text{CNOT}\}$ generates the Clifford group, a flag FTQEC protocol with two-tailed adaptive time decoder or separated $X$ and $Z$ counting is applicable to any fault-tolerant Clifford computation given that the CSS code is self-orthogonal. Note that the magic state distillation and injection [56, 57] use only Clifford operations. Thus, our techniques are also applicable to fault-tolerant universal quantum computation given that high-fidelity magic states are provided.
@@ V.1 Methods
Our optimization tools for space and time decoders including the compact lookup table construction, the MIM technique, and the adaptive time decoders for flag FTQEC are applicable to any stabilizer code. However, we focus on a specific family of codes where the aforementioned tools can be simplified and extended techniques, including separated $X$ and $Z$ decoding and classical processing of flag information are applicable—the family of self-orthogonal CSS codes in which the number of physical qubits is odd, the number of logical qubits is 1, and logical $X$ and $Z$ operators are transversal. To evaluate the performance of our tools, we simulate FTQEC protocols on the $[\![(3d^{2}+1)/4,1,d]\!]$ hexagonal color codes [41] of distance 3, 5, 7 and 9. These codes are planar topological codes with configurations displayed in Fig. 6. For each code, stabilizer generators are measured using the syndrome extraction circuits with single flag ancilla, as depicted in Fig. 4. It was proven that for the hexagonal code of any distance, using flag circuits of this form preserves the code distance regardless of the gate orderings [32, 35]). The simulation is implemented under the circuit-level depolarizing noise model specified in Section II.1. As there is no idling noise in our error model, the syndromes can be extracted sequentially.
@@
FIGURE 6: The studied members of the hexagonal color code, for distances 3, 5, 7, and 9 (right to left). Qubits are on the vertices and stabilizer generators are the plaquettes. As the codes are self-orthogonal CSS codes, both the $X$ and $Z$ stabilizer generators are described by the same layout.
@@ V.2 The overall effect of optimization tools
We first compare two protocols: (1) the FTQEC protocol with Shor time decoder without the MIM technique (the protocol in which none of our optimization tools are applied) and (2) the FTQEC protocol with the MIM technique and the two-tailed adaptive time decoder with the ZX strategy (the best FTQEC protocol in this work which is compatible with any Clifford computation on a self-orthogonal CSS code). The logical error rate $p_{L}$ vs physical error rate $p$ for hexagonal color codes of distance 3, 5, 7, and 9 are plotted in Fig. 7. Our results show that for each code, applying the optimization tools can significantly improve the pseudothreshold (the intersection between each plot and the $p_{L}=2p/3$ line). Furthermore, the optimized decoder yields orders of magnitude improvements in the logical error rate in the $p=10^{-4}$ error regime.
@@ VI Discussions and conclusions
In this work, we focus on flag FTQEC with lookup table decoding and improvements to a decoder consisting of a time decoder and a space decoder. For the space decoder, we first develop a technique to build the lookup table more efficiently in Section III.1. With our lookup table construction method, the lookup table for a self-orthogonal CSS code requires at least 87.5% less memory compared to the lookup table for a generic stabilizer code. The construction method also verifies the distinguishability of the fault set corresponding to flag circuits for syndrome measurements. Our construction also leads to the notion of the fault code, a linear code corresponding to the faults under circuit-level noise, which simplifies the verification of the distance of the protocol. More efficient decoding schemes for the fault code can be an interesting avenue to explore in future work.
@@
For the time decoder, we generalize the adaptive syndrome measurement technique from the previous work [38] (which is applicable to Shor-style error correction [1]) to flag FTQEC, and develop one-tailed and two-tailed adaptive time decoders in Section IV.2. For a general stabilizer code in which flag FTQEC is possible, the one-tailed decoder is preferable as it is compatible with any fault-tolerant quantum computation, while the two-tailed decoder is applicable to quantum memory only. Nevertheless, for self-orthogonal CSS codes, the two-tailed decoder is applicable to any fault-tolerant computation built from Clifford gates and application of $T$ gates by gate teleportation using high-fidelity magic states with the help of the classical processing technique on cumulative flag vectors developed in Section IV.3. The effect of the adaptive time decoders on the simulated storage is discussed in Section V.4. We observe that our adaptive time decoders can improve the pseudothresholds compared to the non-adaptive (Shor) time decoder while preserving the code distance. The two-tailed decoder also outperforms the one-tailed decoder.
@@
We can further improve the performance of adaptive time decoders on self-orthogonal CSS codes by using the separated $X$ and $Z$ counting technique described in Section IV.3. Here, we estimate the number of faults occurred from the measurement of generators of the first type (either $X$ or $Z$) and then use that information in the measurement of generators of the second type. The effect of this technique can be found in Section V.5. When the logical $|0\rangle$ state is stored, we find that the protocol that measures $Z$-type generators before $X$-type generators performs the best. We see no significant difference in the protocol that measures $X$-type generators before $Z$-type generators, and the protocol that measures $X$-type and $Z$-type generators jointly. Thus, the separated $X$ and $Z$ counting provides an advantage only for certain input states depending on the measurement order.
@@ Appendix A Memory footprint savings in the lookup table
If the code is self-orthogonal, then the two tables coincide, $T:=T_{X}=T_{Z}$, $r:=r_{X}=r_{Z}$. Thus,
$M_{CSS,CRO,SO}$ $=\frac{1}{2}\left(2(r_{Z}T_{Z}+r_{X}T_{X})+2n+k(T_{Z}+T_{X})\right)$
@@
which is consistent with having $T+1$ entries, with a map key of $n-k$ bits (with $(n-k)/2$ for flags and for generator bits) and a value with $k$ bits. If we would view the self-orthogonal CSS code as a generic stabilizer code, we would get $T_{stab}=1+2T+T_{XZ}$ and thus the ratio is,
$M_{CSS,CRO,SO}/M_{stab}$ $=\frac{n(T+1)}{T_{stab}(4n-2k)}$
@@
where we used the fact that $T_{XZ}$ must be at least 1 for $t>0$ and a non-trivial encoding. This upper bound means that at a zero rate code, leveraging the structure of a self-orthogonal CSS code and the CROs can create a memory footprint less than 12.5% that of the memory footprint of a lookup table if we viewed the code as a stabilizer code.